OpenStudy (anonymous):

Hi everybody! Find all positive integral solutions (x,y) of the polynomial equation: 4x^3+4x^2y-15xy^2-18y^3-12x^2+6xy+36y^2+5x-10y=0

OpenStudy (anonymous):

any thoughts/ideas?

OpenStudy (angela210793):

couldn't u find a longer eq? :P

OpenStudy (anonymous):

its a monstrous beast...

OpenStudy (angela210793):

:P

OpenStudy (angela210793):

I think it is smth like (ax-by)^3 bu i can't tell wht r a and b

OpenStudy (angela210793):

but*

OpenStudy (anonymous):

good try angela..but u never know that will factorize completely to 3 linear factors!... Infact this BEAST can be factored to any one of the following: 1.3 linears. 2.1 linear and 1 conic

OpenStudy (anonymous):

(x-2*y)*(2*x-1+3*y)*(2*x-5+3*y)

OpenStudy (anonymous):

(x-2y)(2x+3y-5)(2x+3y-1)=0

OpenStudy (angela210793):

i'm not planning to deal with this..u do it...:P

OpenStudy (anonymous):

i got it first, please medal

OpenStudy (anonymous):

u rock cantorset! but we cant use comps!

OpenStudy (anonymous):

how do you know i use computer

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@contorset: lol, its obvious. :P

OpenStudy (angela210793):

hey how did u do tht?

OpenStudy (anonymous):

well witout at computer its LUCK..

OpenStudy (anonymous):

I didn't do it on my own, wolframalpha

OpenStudy (anonymous):

maple

OpenStudy (angela210793):

and how do u do it in comp?

OpenStudy (anonymous):

haha... if u had a good luck u wud put x=2y and inspect!!!

OpenStudy (anonymous):

you can do it in wolframalpha,

OpenStudy (anonymous):

then u get (x-2y) and get others by dividing

OpenStudy (anonymous):

i should not give out my secrets though

OpenStudy (anonymous):

k. but tat doesnt solve the problem..

OpenStudy (anonymous):

x=2y ? you mean substitute

OpenStudy (anonymous):

@saubhik: You need a VERY good luck for that!

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

but see tat doesnt..solve the problem

OpenStudy (anonymous):

we have to find positive integer solutions..

OpenStudy (angela210793):

woww

OpenStudy (anonymous):

but you can pick anything, x = 4y for instance

OpenStudy (angela210793):

and i thought u guys were smart :P:P:P

OpenStudy (anonymous):

i dont see any logic in this problem

OpenStudy (anonymous):

ya cantorset that's truly possible but as i said we need luck and brain-breaking insights...

OpenStudy (anonymous):

see..now how to find the solution?

OpenStudy (anonymous):

for the first one(x-2y), x = 2I y=I

OpenStudy (anonymous):

u can have infinite pairs (2n,n) for a positive integer n since x-2y is a factor...what for the other two?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what for 2x+3y-5=0

OpenStudy (anonymous):

its obvious (1,1) but are there others..THIS IS DIOPHANTINE!!

OpenStudy (anonymous):

OK lets see, x + 3y/2 =5/2 x=(5-3y)/2 5-3y should be even,

OpenStudy (anonymous):

so y is odd.

OpenStudy (anonymous):

what about 2x+3y=1? yay...theres no positive integers here in this equation..so this is infertile:P

OpenStudy (anonymous):

Why do you think so?

OpenStudy (anonymous):

how can u get postive integers (x,y) from 2x+3y<5

OpenStudy (anonymous):

there are none...:P

OpenStudy (anonymous):

infertile..mother has no babies:P

OpenStudy (anonymous):

that can be solved with euclids algorithm

OpenStudy (anonymous):

Oh sorry, i dint read that 'positive' part! :P

OpenStudy (anonymous):

this mother 2x+3y=5 has only one baby i.e. (1,1)

OpenStudy (anonymous):

So our answer: (1,1) and (2n,n)

OpenStudy (angela210793):

OMG!!! This is giving me headache only by reading ur explanations imagine if i had to solve this....:@@@@@@@@

OpenStudy (anonymous):

So its done, (n,2n) U (1,1)

OpenStudy (anonymous):

Openstudy rocks!! thnx evrybody!

OpenStudy (anonymous):

Oops (2n,n). yes right

OpenStudy (angela210793):

Going to deal with literature bye and have a good time :)

OpenStudy (anonymous):

, amog, you wrote x= 1, y = 21?

OpenStudy (anonymous):

@angela: buh-bye. @cantorset: It was an 'I' (eye)

OpenStudy (anonymous):

for the first one(x-2y), x = 2I y=I

OpenStudy (anonymous):

Yes i corrected it, didn't I?

OpenStudy (anonymous):

hi cantorset...do u mind givin me an idea how to have a factor to the eqn..

OpenStudy (anonymous):

suppose i have the baddest luck and go on inspecting from x=3y;x=4y....

OpenStudy (anonymous):

wait, ... start over . why would i think to use x = 3y or anything like that

OpenStudy (anonymous):

also its 3 linear, not a conic

OpenStudy (anonymous):

or worser x=3y+7 etc..

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

you mean because the solution is ( ax + by ) ( dx + cy + ...

OpenStudy (anonymous):

See cantormath..the given equation Infact this BEAST can be factored to any one of the following: 1.3 linears. 2.1 linear and 1 conic

OpenStudy (anonymous):

now clearly theres no constant term in this equation...so a linear factor (which must be present) must pass through the origin...

OpenStudy (anonymous):

and so we start checking by x=y;x=2y;x=3y...

OpenStudy (anonymous):

if the question does not mean to torment us we can get the factor in a few trials...

OpenStudy (anonymous):

k..thnx evrybody for helping..:)

OpenStudy (anonymous):

wait, youre saying in general, a cubic in x and y can be factored into either 3 linears, or 1 linear and 2 conics? , or 2 linears, and a conic?

OpenStudy (anonymous):

what do you mean by constant term in the equation

OpenStudy (anonymous):

term with no variables attched..

OpenStudy (anonymous):

oh right

OpenStudy (anonymous):

the final solution is 3 linears, right?

OpenStudy (anonymous):

yups...family of 3 straight line

OpenStudy (anonymous):

cool, so ok , theres no constant term, and clearly, (0,0) works

OpenStudy (anonymous):

yups..try looking for x=y,2y,3y,...here u need luck

OpenStudy (anonymous):

so that tells us that x = ay or y = ax must be one of the solutions?

OpenStudy (anonymous):

yes..true

OpenStudy (anonymous):

but only for a=2

OpenStudy (anonymous):

i.e.(2n,n)

OpenStudy (anonymous):

right, but thats only if you do the long division< right?

OpenStudy (anonymous):

if you know to assume, 2x = y ?

OpenStudy (anonymous):

or wait, you know that (0,0) is a solution and (1,1) is a solution, right>

OpenStudy (anonymous):

yes...not assume but see if its correct by replacing 2x by y OR replacing y by 2x

OpenStudy (anonymous):

(1,1) is known after we get other factors using x-2y

OpenStudy (anonymous):

hmmm, yes but i mean is there something that would make us want to to do x = 2y

OpenStudy (anonymous):

oh ,

OpenStudy (anonymous):

tat is LUCK!

OpenStudy (anonymous):

( 1,-1) is another solution then

OpenStudy (anonymous):

OR if u knew the problem beforehand OR if u have wolframalpha open on ur browser. No...-1 is not a POSITIVE integer :P

OpenStudy (anonymous):

4x^2 + 9y^2 + 12xy -2x -15y +5 =0 How do you solve this btw?

OpenStudy (anonymous):

from discriminant criteria for various conics see if its parabola/ellipse/hyperbola..

OpenStudy (anonymous):

k. for positive integer solutions...ALL 2-VARIABLE EQUATION DO NOT HAVE positive integer solutions.....such questions are given after testing them..